A piggy bank contains 20 coins in 5 centavos and 1 peso. Their total value is 8.75 pesos. How many of each kind are present?
Answer by Boreal(15235)

This can't be solved. The centavos have to be 15 in order to get 0.75
The pesos then are 5, and the cost is 5.75.
Am I missing something? a centavo is 1/20th of a peso, right?
x=5 centavo number
20-x= peso
x*(.05) + (20-x)1=8.75 This relates the number of coins to the cost.
.05x + 20-x = 8.75
collect terms and subtract 20 from both sides
-0.95 x = -11.25
Multiply both sides by (-1)
0.95 x = 11.25
This is not an integer. It is close to 12 but NOT 12.
12 centavos = 0.60
8 pesos =8.00 Doesn't work
15 centavos=0.75
5 pesos=5.00 Doesn't work
No solution!!!